# A Survey on Wi-Fi Sensing Generalizability: Taxonomy, Techniques, Datasets, and Future Research Prospects

### v1

@@ 3. Dataset
34 Person-in-WiFi 3D (Yan et al., 2024b) 3 environment, 7 participants, 8 actions 20MB,5GHz https://aiotgroup.github.io/Person-in-WiFi-3D/
35 $M$-Fi 4 environment, 40 participants, 27 actions 20MB,5GHz https://ntu-aiot-lab.github.io/$m$-fi
@@
3 ARIL (Wang et al., 2019a) 1 environment, 16 locations, 1 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/geekfeiw/ARIL
36 Chen and Chang 2022 (Chen and Chang, 2022) 2 environment, 7 participants, 16 actions 20MB,5GHz https://github.com/BingJiaChen/few-shot-indoor-localization
37 MetaLoc (Gao et al., 2023) 2 environment, 180 locations 20MB,5GHz https://github.com/StatFusion/MetaLoc
@@
38 DASECount (Hou et al., 2022b) 16 locations, 1 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz http://bisuzhi.cn/Sensedata.html
@@
39 XRFv2 (Lan et al., 2025) 3 environment, 16 participants, 31 actions 1Tx, 3Rx, 20MB, 5GHz https://github.com/aiotgroup/XRFV2
Table 1. Open-sourced Dataset in Wi-Fi Human Sensing. (Continued.)
  1. (1)
@@
    XRF55 (Wang et al., 2024b) dataset was collected in four indoor environments. In each environment, users performed 55 types of actions, categorized into five categories: human-object interactions, human-human interactions, fitness, body motions, and human-computer interactions, within a 3.1m $\times$× 3.1m area. The first scene involved 30 participants, while the other three scenes involved 3 participants each. Each participant performed each action 20 times, with 42,900 samples in total, lasting 59.58 hours. The data collection setup included one transmitter with a single antenna working, and three receivers each with three antennas, arranged in a quadrilateral layout. The sampling rate was 200 Hz. Notably, XRF55 dataset also includes synchronized modalities from millimeter-wave radar, RFID, and RGB+D+IR, supporting multimodal learning or cross-modal tasks.
  2. (2)
    Person-in-WiFi 3D (Yan et al., 2024b) dataset was collected in three environments: an office, a classroom, and a corridor, each within a 4m $\times$× 3.5m area. Seven participants freely performed eight actions: reaching out, raising hands, bending over, stretching, sitting down, lifting legs, standing, and walking. In addition to scenarios with a single participant, there were also cases where 2, 3, or 4 participants simultaneously performed actions within the scene. Each clip was recorded for 40 seconds, resulting in a total of 456 40-second recordings. The data collection setup included one transmitter with a single antenna and three receivers, each with three antennas, arranged in a quadrilateral layout. The sampling rate was 200 Hz.
  3. (3)
    ARIL (Wang et al., 2019a) dataset was collected in a 3.2m $\times$× 3.2m area within a laboratory, with 16 positions. The actions include 6 hand gestures for potential human-computer interaction applications: hand up, hand down, hand left, hand right, hand circle, and hand cross. One volunteer performed each gesture 15 times at each position. The dataset contains a total of 1,394 valid samples. Data was collected using two USRPx210 devices, each with a single antenna, as transmitter and receiver. The sampling rate was approximately 60Hz.
  4. (4)
@@
    BullyDetect (Lan et al., 2024) dataset was collected in 8 different environments: a classroom, corridors, playgrounds, rooftops, a meeting room, a corner of the campus, an Office, and an abandoned building. Five volunteers were paired in 10 pairs, each acting as either an attacker or a victim during the experiment. The participants performed 7 different bullying actions: pushing, kicking, slapping, grabbing, punching, kneeing, and hitting with a stick. Each action was performed 5 seconds and 10 times, resulting in 11,200 samples. The experimental setup included a transmitter with a single antenna and one 3-antenna receiver, spaced 3 meters apart, with a sampling rate of 1000 Hz.
  5. (5)
[…]
  6. (6)
    GaitID (Zhang et al., 2020c) dataset is designed for gait recognition tasks. Data was collected in two scenarios: a hall and a discussion room. Volunteers walked along four predefined tracks, starting from both ends of each track. In the hall, 10 volunteers participated, with each walking 50 times in each direction. In the discussion room, 3 volunteers repeated each track 25 times in each direction. A total of 4,600 samples were collected. The data collection setup included one transmitter and six receivers. The sampling rate was not reporeted.
  7. (7)
    WiAR (Guo et al., 2019) dataset was collected in three indoor environments: one empty room, one meeting room, and one office, as well as two outdoor environments, including a playground and a corridor. Ten volunteers performed three categories of actions: upper body activities, lower body activities, and whole-body activities, with 16 different actions. Each action was performed 30 times. The experimental setup included a transmitter with a single antenna and one receiver with three antennas, spaced 4 meters apart. The participants performed the actions at the center of the setup. The sampling rate was 30Hz.
  8. (8)
    SignFi (Ma et al., 2018) dataset was collected in two environments: a laboratory and a home. The lab, measuring 13m $\times$× 12m, the distance between the AP and the STA was 230 centimeters. While the home environment, with dimensions of 4.11m $\times$× 3.86m, offered a shorter distance of 130 centimeters between the AP and STA. 5,520 instances were recorded in the lab, while the home environment provided 2,760 instances. The data involved 276 different sign language gestures, encompassing movements of the head, arm, hand, and fingers. The gestures were performed by 5 different users, and the users were instructed to perform each sign gesture repeatedly, with 20 instances for the lab environment and 10 instances for the home environment, resulting in a total of 8280 gesture instances.The sampling rate was 200Hz.
  9. (9)
    Meneghello et.al. 2023 (Meneghello et al., 2023) dataset was collected across seven diverse environments including a bedroom, living room, kitchen, university laboratory, university office, semi-anechoic chamber, and a meeting room. Thirteen participants consisting of ten males and three females freely performed eight actions: walking, running, jumping, sitting still, standing still, sitting downs, standing up, and arm exercises. And the dataset consists of more than 13 hours of data (amounting to 23.6 GB of raw CSI readings).
  10. (10)
[…]
  11. (11)
    Yousef et.al. 2017 (Yousefi et al., 2017)dataset was collected in an indoor office area where the WiFi transmitter and receiver were positioned 3m apart in a line-of-sight (LOS) condition. Six volunteers performed six different activities: lying down, falling down, walking, running, sitting down, and standing up. Each activity was performed 20 times, and each activity was recorded for 20 seconds. Additionally, video recordings were made during the experiments to help label and annotate the activities in the dataset.The sampling rate was 1 kHz.
  12. (12)
@@
    WiSDA (Jiao et al., 2024) dataset was collected in an office setting, equipped with a desk and chair, within a $\times×2m$ sensing area. Five participants performed six activity categories: Push, Sweep, Tap, Slide, Draw-Z, and Draw-N. Each person collected 100 samples on 30 subcarriers and converted them to images, resulting in 3000 samples per action. The data was captured using a single pair of antennas, with one acting as the transmitter and the other as the receiver, positioned to face each other within the sensing area, similar to the setup used in the Widar 3.0 dataset.
  13. (13)
[…]
  14. (14)
    CSI-HAR-Dataset (Moshiri et al., 2021) dataset was collected in three environments incluing a classroom, a library, and a park. The dataset was collected by 6 volunteers including three male and three female. Each volunteer performed five categories of actions: sitting, standing, walking, running, and stretching. Sitting involved participants holding different postures on a chair for 5 seconds, while standing required them to maintain different body orientations for 5 seconds. Walking consisted of participants covering a distance of 10m at a normal pace, and running involved moving 10m at a moderate pace. Stretching included bending down, stretching arms overhead and so on. And each action was repeated 20 times by each participant. The experimental setup involved two cameras positioned 3m apart in the indoor spaces and 5m apart in the park. Additionally, a motion sensor was placed at the center of the test area, 4m away from the participants during indoor trials and 7m for the outdoor park trials. The sampling rate was 60Hz.
  15. (15)
@@
    WiMANS (Huang et al., 2024) dataset was collected in three environments including a classroom, a meeting room and an empty room. Each environment was equipped with a transmitter, a receiver and a monitor camera used to capture synchronized videos. The activities were performed by five volunteers including an adult, a middle-aged person, and an elderly person. Nine activities were included in the dataset: nothing, walking, rotation, jumping, waving, lying down, picking up, sitting down and standing up. Each environment has five different locations. The dataset provides 11286 samples (over 9.4 hours) of dual-band WiFi CSI and synchronized videos.
  16. (16)
    ImgFi (Zhang and Jiao, 2023) dataset was collected in a laboratory, where 5 volunteers performed 6 actions: bending, drinking, nodding, squatting, drawing the letter O, and drawing the letter B. Each action was repeated 20 times, resulting in a total of 600 samples. Data collection was conducted using one transmitter and one receiver positioned 4 meters apart, with the volunteers performing actions in the center. The sampling rate was 100Hz.
  17. (17)
    ReWis (Bahadori et al., 2022) dataset was collected in three environments: an office, a meeting room, and a classroom. Two participants performed three actions—jumping, walking, and standing—10 times in each environment, with each session lasting 180 seconds and a minimum interval of 2 hours between sessions. A Netgear R7800 Wi-Fi router with a Qualcomm Atheros chipset operated in AP mode, while an off-the-shelf laptop functioned as the client. Additionally, three Asus RT-AC86U Wi-Fi routers, each equipped with four antennas, recorded Wi-Fi CSI at a rate of 100 Hz.
  18. (18)
    FallDeFi (Palipana et al., 2018), designed for fall detection, was collected in five environments: a kitchen, bedroom, corridor, laboratory, and bathroom. Three participants performed fall actions and other activities such as walking, jumping, sitting down, and standing up, with each action lasting for 10 seconds and 1153 samples in total. Each environment featured one transmitter and one receiver, with varying placement distances between 5 and 9 meters depending on the setting. To introduce environmental variations, data was collected on different days, and the transmitter’s position was adjusted during the experiments. All experiments were conducted at 5.2 GHz with a sampling rate of 1000 Hz.
  19. (19)
    FallDar (Yang et al., 2022b) dataset, developed for fall detection, was collected in two environments: a home environment and an office environment. In the home environment, five participants performed fall actions at three locations—dining room, living room, and balcony—resulting in 115 fall samples. The transmitter and receiver were placed in two settings: line-of-sight (LoS) and non-line-of-sight (NLoS). In the office environment, the transmitter was fixed in position, while the receiver was placed at 16 locations sequentially. Two participants contributed to the dataset. One performed fall actions at locations 1 and 3, while the other performed falls at locations 1 and 4. Additionally, one participant walked freely in the office to collect normal activity data. This resulted in 228 fall samples and 504 normal activity samples. In both environments, the transmitter had one antenna, and the receiver had three antennas. Data was collected at a sampling rate of 1,000 Hz with a center frequency of 5.825 GHz.
  20. (20)
    SAR  (Brinke and Meratnia, 2019a, c) dataset was collected in the living room of an apartment. A total of 9 volunteers participated in experiments over 6 days. During the first 3 days, 3 different volunteers participated each day, while 2 volunteers participated repeatedly over the last 3 days. During the first 3 days, participants performed 6 actions: Clap, Walk, Wave, Jump, Sit, and Fall. In the last 3 days, participants performed the same actions except for Jump, resulting in a total of 5 actions. Each action was performed continuously for 5 seconds and repeated 50 times. The data collection setup consisted of a transmitter and a receiver, each equipped with three antennas, spaced approximately 2.5 meters apart. The sampling rate was 50 Hz.
  21. (21)
@@
    WiGuesture (Zhao et al., 2024b) dataset was collected in a conference room. Eight volunteers participated in the study. The participants performed six distinct gestures: moving left-right, forward-backward, up-down, circling clockwise, clapping, and waving. Each gesture was recorded continuously for one minute. The data collection setup consisted of a six-antenna home Wi-Fi router acting as the transmitter and a single-antenna ESP32S3 microcontroller serving as the receiver. The transmitter and receiver were placed 1.5 meters apart, and data was captured at a sampling rate of 100 Hz.
  22. (22)
    DeepSeg  (Xiao et al., 2020) dataset was collected in a meeting room with data from 5 volunteers performing 10 activities. These actions include 5 fine-grained movements such as swinging the hand, raising the hand, making a pushing motion, tracing a circular pattern (O), and tracing a cross pattern (X), as well as 5 coarse-grained activities, namely boxing, picking up an object, running, squatting down, and walking. Each participant repeated each action 30 times, resulting in 1,500 samples. The dataset collection setup included a single-antenna transmitter and a three-antenna receiver placed 2 meters apart, with a sampling rate of 50 Hz.
  23. (23)
    OPERAnet (Bocus et al., 2022) dataset was collected in two rooms and includes data for seven activities: walk, sit, stand from a chair, lie down, stand up from the floor, rotate the upper-half body, and steady state, totaling 417 minutes of activity data. Additionally, the dataset includes 18 minutes of background data and 27 minutes of crowd-counting data. The data collection setup consisted of one transmitter with three antennas and two receivers, each with three antennas, arranged in a right-angled triangular configuration. The second receiver was positioned at the right-angle vertex of the triangle. The sampling rate for Wi-Fi signals was 1600 Hz. OPERAnet dataset also includes synchronized data from two Kinect cameras, four UWB nodes, and a Passive WiFi Radar system constructed using a USRP-2945. This multimodal setup supports research in both multi-modal learning and cross-modal tasks.
  24. (24)
    Chen and Chang (Chen and Chang, 2022) dataset is used for indoor positioning based on CSI (Channel State Information) fingerprints and includes two environments: conference and office. In the conference scene, there is one pair of transmitters and receivers, and data is collected at 16 different locations, with 600 CSI fingerprint samples recorded at each location, resulting in a total of 9600 samples. In the office scene, two pairs of transmitters and receivers are used, with data collected at 18 locations, and 500 CSI fingerprint samples recorded at each location, yielding a total of 9000 samples.
  25. (25)
    HTHI (Alazrai et al., 2020) dataset was collected in a furnished room. 66 volunteers participated in the study, forming 40 unique pairs to perform 12 types of human-to-human interaction activities, such as handshaking, high fives, and hugging. Each pair performed each activity 10 times, resulting in 4,800 samples. The data collection setup consisted of a commercial off-the-shelf Sagemcom 2704 router with two antennas acting as the transmitter and a desktop computer with three antennas serving as the receiver. The transmitter and receiver were positioned 4.3 meters apart. The sampling rate for the data collection was not reported.
  26. (26)
    MM-Fi (Yang et al., 2024a) dataset comprises data collected in two rooms, 4 environmental settings, with 10 participants in each environment. The participants performed 14 daily activities, such as raising the left arm, waving the left arm, and waving the right arm, as well as 13 rehabilitation exercises, such as left front lunge and right front lunge. The setup included a transmitter with one antenna and a receiver with three antennas, arranged with a spacing of 3.75 meters between the receiver antennas. Participants performed actions while positioned 3 meters away from the receiver, facing the receiver. Each action was performed for 30 seconds. The devices operated at 5 GHz with a sampling rate of 1000 Hz. In addition to Wi-Fi sensing, MM-Fi includes data from LiDAR, mmWave radar, and camera modalities, supporting multimodal learning tasks.
  27. (27)
[…]
  28. (28)
    NTU-Fi HAR (Yang et al., 2022a) was collected in a laboratory setting with 20 participants performing six types of actions, including running, walking, falling down, boxing, circling arms, and cleaning the floor. Each participant performed each action 20 times, resulting in a total of 2,400 samples. The setup consisted of a transmitter with one antenna and a receiver with three antennas. The equipment operated at 5 GHz with a sampling rate of 500 Hz.
  29. (29)
    CrossSense (Zhang et al., 2018) dataset was collected in three scenarios: a hall entrance, a narrow corridor, and a room, with the participation of 100 volunteers. For gait recognition, each volunteer walked 20 times in the same direction in each scenario, resulting in a total of 6,000 samples. For action recognition, five predefined positions were set in each scenario. At each position, each volunteer performed 40 different actions (e.g., pull, drag, throw), repeating each action 10 times. This resulted in a total of 600,000 samples. The experimental setup included one transmitter and one receiver, both equipped with three antennas. The devices operated at 5 GHz with a sampling rate of 1,000 Hz. Currently, access to the CrossSense dataset requires completing a consent form in its github repository.
  30. (30)
    RF-Net (Ding et al., 2020a) dataset was collected in six different rooms with 11 volunteers performing six types of actions: wiping, walking, moving, rotating, sitting, and standing up. The Wi-Fi setup included two transmitter-receiver pairs, operating at a sampling rate of 100 Hz.
  31. (31)
    CSIDA, related in WiGr (Zhang et al., 2021c), was collected in an office environment with three distinct locations and a classroom environment with two locations. Five volunteers performed six gestures: pull left, pull right, lift up, press down, draw a circle, and draw zigzag. Each gesture was repeated 10 times, with each repetition lasting 1.8 seconds. The device setup consisted of a single-antenna transmitter and a three-antenna receiver, separated by a distance of 2.6 meters. The system operated in the 5 GHz mode with a sampling rate of 1000 Hz.
  32. (32)
    RISE (Zhai et al., 2021) dataset includes data for both gesture recognition and gait recognition, collected in two environments: an office and a controlled setting featuring an RF anechoic chamber to minimize multipath effects. Gesture Recognition: six volunteers performed six gestures—push and pull, draw a circle, throw, slide, sweep, and draw zigzag—across five locations in both environments. Each participant repeated each gesture 30 times at each location, resulting in a total of 10,800 samples. Gait Recognition: 15 volunteers walked under five different configurations in the controlled environment. Each volunteer completed 30 walks, generating 1,150 samples in total. The setup consisted of one transmitter and one receiver, operating at a sampling rate of 1,000 Hz.
  33. (33)
@@
    RoGER (Bi et al., 2024) dataset was collected in a meeting room at two distinct locations. Location 1: A single volunteer performed six gestures, i.e., Push & Pull, Draw Zigzag, Clap, Sweep, Draw Circle, Slide, the same as Widar 3.0 (Zheng et al., 2019), in four different directions. Each gesture was repeated 30 times, resulting in 480 samples. Location 2: Four volunteers (including the participant from Location 1) performed the same six gestures in four directions. Each volunteer repeated each gesture 30 times, generating 1,920 samples. Each sample lasts for 10 seconds. The setup included a two-antenna transmitter and a three-antenna receiver, operating at 2.437 GHz with a sampling rate of 100 Hz.
  34. (34)
@@
    ResMon (Zheng et al., 2023) dataset was collected in two environments: a laboratory and a meeting room, with the laboratory further collected at two areas. Six volunteers performed 4 actions: stable breath, cough, sneeze, and yawn. The experimental setup included a two-antenna transmitter and a three-antenna receiver, operating at 2.437 GHz with a sampling rate of 100 Hz.
  35. (35)
    DASECount  (Hou et al., 2022b) dataset, designed for crowd counting, was collected in two environments: an office and a lecture hall, each with both LOS and NLOS configurations, resulting four different environmental settings. Volunteers, ranging from 0 to 8 people, performed three types of motion: static, dynamic, and mixed. Each sampling sequence was recorded for 5 minutes. The experimental setup included a two-antenna transmitter and a three-antenna receiver, operating at 2.437 GHz with a sampling rate of 100 Hz.
  36. (36)
    Dataset3 (bis, 2024) is used for fatigue detection and includes two environments: a meeting room and an office room. It contains four types of fatigue actions: yawning, rubbing eyes, nodding, and stillness. Six volunteers performed each action 50 times in the meeting room and 40 times in the office room, resulting in a total of 2160 samples from both environments.
  37. (37)
    MetaLoc (Gao et al., 2023) dataset was designed for indoor localization and collected in a 12 m $\times$× 5 m hall and a 10 m $\times$× 8 m laboratory. In each environment, Wi-Fi fingerprints were recorded at 90 locations. The setup included three transmitters: an ASUS RT-AC86U, a TP-Link TL-WR885N, and a TP-Link TL-WR886N, and a Nexus-5 smartphone as the receiver. The system operated at 5 GHz.
  38. (38)
    Baha’A et.al. 2020 (Baha’A et al., 2020) data collected in three different environments: two with line-of-sight (LOS) settings and one with a non-line-of-sight (NLOS) setting. Thirty volunteers performed five different actions: falling from a sitting position, falling from a standing position, walking, sitting down and standing up, and picking up a pen from the ground. Each action was performed 20 times. The experimental setup included a single-antenna transmitter and a three-antenna receiver, operating at a frequency of 2.4 GHz with a sampling rate of 320Hz.
  39. (39)

### v2

@@ 31 CrossSense [158] 3 environments, 100 participants 1Tx, 1Rx, 20MHz, 5GHz 100 participants
35 XRFv2 [296] 3 environments, 16 participants, 30 actions 1Tx, 3Rx, 20MHz, 5GHz Multimodal, continuous actions
36 Person-in-WiFi 3D [50] 3 environments, 7 participants, 8actions 1Tx, 3Rx, 20MHz, 5GHz Multi-person
1 XRF55 [70] 4 environments, 39 participants, 55 actions 1Tx, 3Rx, 20MHz, 5GHz Multimodal, 55 actions
12 MM-Fi [72] 4 environments, 40 participants, 27 actions 20MHz, 5GHz Multimodal
@@
37 MetaLoc [197] 2 environments, 180 locations 3Tx, 1Rx, 20MHz, 5GHz RSSI and CSI
38 Chen and Chang [297] 2 environments, 16 locations 2Tx, 2Rx CSI fingerprints
26 ARIL [15] 1 environments, 16 locations, 1 participants, 6 actions 1Tx, 1Rx, 20MHz, 2.4GHz USRP data, clean phase
@@ 38 Chen and Chang [297] 2 environments, 16 locations 2Tx, 2Rx CSI fingerprints
39 SDP [298] 4 scenarios multi-APs, 80/160MHz, 5GHz 0-2 people
40 DASECount [227] 2 environments, 2 scenarios (NLOS/LOS), 3 motions 1Tx, 1Rx, 40MHz, 2.4GHz 0-8 people
10 Meneghello et al. [286] 1 environments, 10 participants 1Tx, 1Rx, 80MHz, 5.21GHz 1-10 people
@@ 39 SDP [298] 4 scenarios multi-APs, 80/160MHz, 5GHz 0-2 people
35 XRFv2 [296] 3 environments, 16 participants, 30 actions 1Tx, 3Rx, 20MHz, 5GHz Multimodal, continuous actions
41 WiFiTAD [299] 1 environment, 3 participants, 7 actions 1Tx, 1Rx, 5GHz Continuous actions
Multiple Sub-Datasets
42 CSI-Bench [73] 26 environments, 35 participants 16 device configuration In-the-wild data, 461 hours
10 Meneghello et al. [286] 7 environments, 13 participants 1Tx, 1Rx, 80MHz, 5.21GHz Three tasks, 80MHz
43 CSI-Net [210] 1 environments, 5 positions, 30 participants 1Tx, 1Rx, 20MHz, 5GHz Four tasks, Biometrics
Figure 17: Bubble chart of public Wi-Fi sensing datasets, where the bubble size indicates the number of samples and the vertical axis represents the number of classes. For visual clarity, datasets with extreme values, such as SignFi [14] (296 classes) and CSI-Bench [73] (over 1,000,000 samples), are excluded from the plot. Similarly, datasets like WiFiTAD [299], MM-Fi [72], and XRFv2 [296], which collect continuous sequences over time without clear per-sample annotations, are also not included due to the ambiguity in estimating their sample counts.
@@ 42 CSI-Bench [73] 26 environments, 35 participants 16 device configuration In-the-wild data, 461 hours
(1) XRF55 [70] dataset was collected in four indoor environments. In each environment, users performed 55 types of actions, categorized into five categories: human-object interactions, human-human interactions, fitness, body motions, and human-computer interactions, within a 3.1m $\times$ 3.1m area. The first scene involved 30 participants, while the other three scenes involved 3 participants each. Each participant performed each action 20 times, with 42,900 samples in total, lasting 59.58 hours. The data collection setup included one transmitter with a single antenna working, and three receivers each with three antennas, arranged in a quadrilateral layout. The sampling rate was 200 Hz. Notably, XRF55 dataset also includes synchronized modalities from millimeter-wave radar, RFID, and RGB+D+IR, supporting multimodal learning and cross-modal tasks. Additionally, XRF55 also releases 2D human pose annotations for the pose estimation task.
@@
(2) BullyDetect [211] dataset was collected in 8 different environments: a classroom, corridors, playgrounds, rooftops, a meeting room, a corner of the campus, an Office, and an abandoned building. Five volunteers were paired in 10 pairs, each acting as either an attacker or a victim during the experiment. The participants performed 7 different bullying actions: pushing, kicking, slapping, grabbing, punching, kneeing, and hitting with a stick. Each action was performed 5 seconds and 10 times, resulting in 11,200 samples. The experimental setup included a transmitter with a single antenna and one 3-antenna receiver, spaced 3 meters apart, with a sampling rate of 1000 Hz.
@@
(3) WiSDA [148] dataset was collected in an office setting, equipped with a desk and chair, within a 2m×$\times$2m sensing area. Three participants performed six activity categories: Push, Sweep, Clap, Slide, Draw-Z, and Draw-N. Each person collected 100 samples on 30 subcarriers and converted them to images, resulting in 3000 samples per action. The data was captured using a single pair of antennas, with one acting as the transmitter and the other as the receiver within the sensing area, similar to the setup used in the Widar 3.0 dataset. The bandwidth is 20 MHz.
(4) WiNDR [63] dataset was collected in an indoor office room measuring 4.8m×$\times$3.1m. Three volunteers incluing two males and one female, aged between 25 and 30 years old and weighing between 50 to 65 kg, were asked to perform five distinct hand gestures: drawing a circle (start/accept), crossing hands (stop), clapping (switch context/menu), raising hands (increase/turn up), and lowering hands (decrease/turn down). The volunteers were instructed to perform these gestures in 24 different directions, marked by gray lines on the floor, with each direction spaced 15 degrees apart. These gestures were performed 40 times in each of the 24 directions, sampled every 15 degrees around the subjects. The collection of gesture data was facilitated by two computers, the antennas were extended with 3.5m cables and positioned 1.2m above the ground using tripods.The sampling rate was 200Hz.
@@
(5) WiMANS [283] dataset was collected in three environments including a classroom, a meeting room and an empty room. Each environment was equipped with a transmitter, a receiver and a monitor camera used to capture synchronized videos. The activities were performed by five volunteers including an adult, a middle-aged person, and an elderly person. Nine activities were included in the dataset: nothing, walking, rotation, jumping, waving, lying down, picking up, sitting down and standing up. Each environment has five different locations. The dataset provides 11286 samples (over 9.4 hours) of dual-band WiFi CSI and synchronized videos.
@@
(6) WiGuesture [284] dataset was collected in a conference room. Eight volunteers participated in the study. The participants performed six distinct gestures: moving left-right, forward-backward, up-down, circling clockwise, clapping, and waving. Each gesture was recorded continuously for one minute. The data collection setup consisted of a six-antenna home Wi-Fi router acting as the transmitter and a single-antenna ESP32S3 microcontroller serving as the receiver. The transmitter and receiver were placed 1.5 meters apart, and data was captured at a sampling rate of 100 Hz.
@@
(7) RoGER [225] dataset was collected in a meeting room at two distinct locations. Location 1: A single volunteer performed six gestures, i.e., Push & Pull, Draw Zigzag, Clap, Sweep, Draw Circle, Slide, the same as Widar 3.0 [67], in four different directions. Each gesture was repeated 30 times, resulting in 480 samples. Location 2: Four volunteers (including the participant from Location 1) performed the same six gestures in four directions. Each volunteer repeated each gesture 30 times, generating 1,920 samples. Each sample lasts for 10 seconds. The setup included a two-antenna transmitter and a three-antenna receiver, operating at 2.437 GHz with a sampling rate of 100 Hz.
(8) ImgFi [285] dataset was collected in a laboratory, where 5 volunteers performed 6 actions: bending, drinking, nodding, squatting, drawing the letter O, and drawing the letter B. Each action was repeated 20 times. Data collection was conducted using one transmitter and one receiver positioned 4 meters apart, with the volunteers performing actions in the center. The sampling rate was 100Hz.
@@
(9) ResMon [224] dataset was collected in two environments: a laboratory and a meeting room, with the laboratory further collected at two areas. Six volunteers performed 4 actions: stable breath, cough, sneeze, and yawn. The experimental setup included a two-antenna transmitter and a three-antenna receiver, operating at 2.437 GHz with a 20MHz bandwidth and sampling rate of 100 Hz. ResMon also includes samples collected at 5GHz with 40MHz bandwidth for cross-band evaluation.
(10) Meneghello et al. [286] present a Wi-Fi sensing dataset collected using the 802.11ac protocol with an 80 MHz bandwidth centered at 5,210 MHz. It supports action recognition (AR), person identification (PI), and people counting (PC) tasks. For AR, four participants performed up to seven activities (e.g., walking, running, jumping, sitting still, standing still, sitting down/standing up, arm exercises) in six environments: including a bedroom, living room, kitchen, university laboratory, university office, and semi-anechoic chamber. For PI, 10 individuals were recorded moving freely in a meeting room (7 × 7.5 × 3.5 m). For PC, 1 to 10 people moved simultaneously in the meeting room. The dataset spans over 13 hours of channel state information (CSI) recordings, collected using the Nexmon tool with Netgear X4S AC2600, Asus RT-AC86U, and TP-Link AD7200 routers.
(11) Demrozi et al. [287] dataset was collected in two office environments. The first office, measuring 12m×$\times$6m×$\times$3m, is divided into two blocks of workspaces, each equipped with a Raspberry Pi device and an access point (AP). The second office is smaller, 6m×$\times$4m×$\times$2.75m, and also contains two Raspberry Pi devices and two APs. The Raspberry Pis and APs were paired to collect Channel State Information (CSI). Six participants, including two females and four males, performed five types of activities: entering the office, walking, standing, sitting, and leaving the office. The devices were arranged such that two wireless routers were placed 5 meters apart at a height of 140 cm, creating two separate 2.4 GHz WiFi networks. The total size of the dataset is approximately 70 GB.
(12) MM-Fi [72] dataset comprises data collected in two rooms, 4 environmental settings, with 10 participants in each environment. The participants performed 14 daily activities, such as raising the left arm, waving the left arm, and waving the right arm, as well as 13 rehabilitation exercises, such as left front lunge and right front lunge. The setup included a transmitter with one antenna and a receiver with three antennas, arranged with a spacing of 3.75 meters between the receiver antennas. Participants performed actions while positioned 3 meters away from the receiver, facing the receiver. Each action was performed for 30 seconds. The devices operated at 5 GHz with a sampling rate of 1000 Hz. In addition to Wi-Fi sensing, MM-Fi includes data from LiDAR, mmWave radar, and camera modalities, supporting multimodal learning tasks.
(13) NTU-Fi HAR [288] was collected in a laboratory setting with 20 participants performing six types of actions, including running, walking, falling down, boxing, circling arms, and cleaning the floor. Each participant performed each action 20 times, resulting in a total of 2,400 samples. The setup consisted of a transmitter with one antenna and a receiver with three antennas. The equipment operated at 5 GHz with a sampling rate of 500 Hz.
(14) FallDar [289] dataset, developed for fall detection, was collected in two environments: a home environment and an office environment. In the home environment, five participants performed fall actions at three locations—dining room, living room, and balcony—resulting in 115 fall samples. The transmitter and receiver were placed in two settings: line-of-sight (LoS) and non-line-of-sight (NLoS). In the office environment, the transmitter was fixed in position, while the receiver was placed at 16 locations sequentially. Two participants contributed to the dataset. One performed fall actions at locations 1 and 3, while the other performed falls at locations 1 and 4. Additionally, one participant walked freely in the office to collect normal activity data. This resulted in 228 fall samples and 504 normal activity samples. In both environments, the transmitter had one antenna, and the receiver had three antennas. Data was collected at a sampling rate of 1,000 Hz with a center frequency of 5.825 GHz.
(15) ReWiS [66] dataset was collected in three environments: an office, a meeting room, and a classroom. Two participants performed four actions—empty room, jumping, walking, and standing—10 times in each environment, with each session lasting 180 seconds and a minimum interval of 2 hours between sessions. A Netgear R7800 Wi-Fi router with a Qualcomm Atheros chipset operated in AP mode, while an off-the-shelf laptop functioned as the client. Additionally, three Asus RT-AC86U Wi-Fi routers, each equipped with four antennas, recorded Wi-Fi CSI at a rate of 100 Hz over 20 MHz (52 subcarriers) and 80 MHz (242 subcarriers) bandwidths at 5 GHz.
(16) OPERAnet [290] dataset was collected in two rooms and includes data for seven activities: walk, sit, stand from a chair, lie down, stand up from the floor, rotate the upper-half body, and steady state, totaling 417 minutes of activity data. Additionally, the dataset includes 18 minutes of background data and 27 minutes of crowd-counting data. The data collection setup consisted of one transmitter with three antennas and two receivers, each with three antennas, arranged in a right-angled triangular configuration. The second receiver was positioned at the right-angle vertex of the triangle. The sampling rate for Wi-Fi signals was 1600 Hz. OPERAnet dataset also includes synchronized data from two Kinect cameras, two UWB systems (4 nodes in system 1, 5 nodes in system 2), and a Passive WiFi Radar system constructed using a USRP-2945. This multimodal setup supports research in both multi-modal learning and cross-modal tasks.